OpenStudy (anonymous):

If you ever watch track and field at the Olympics, you'll notice that the starting points of the runners are not even if the race is run on a curve and the runners have to run in lanes. This is to offset the fact that the distances along each lane are unequal on a curve. To see this effect, consider two runners, each of whom runs at a speed of 7 m/s. The runners run on a circular track. The radius of the inside lane is 50 m, and the radius of the outside lane is 51 m. By how many seconds will the inside runner beat the outside runner if they each run once around the track?

OpenStudy (agent0smith):

First find the distance each runner covers - it will be the circumference of a circle. Circumference is 2*pi*r, so the inside track has a circumference of: \[C _{1}=2 \times \pi \times 50\]outside track has a circumference of \[C _{2}=2 \times \pi \times 51\] Then use time = distance/speed for each runner.

OpenStudy (anonymous):

Huhh guide I still confused

OpenStudy (anonymous):

i barely understand the reading

OpenStudy (anonymous):

what is it asking?

OpenStudy (agent0smith):

"By how many seconds will the inside runner beat the outside runner if they each run once around the track?" It's asking if they do one lap, ie one circumference of the circle, what is the difference in their times?

OpenStudy (anonymous):

ohh... sorry

OpenStudy (anonymous):